OpenStudy (anonymous):

How to solve 2e^ln x + 3 = 6?

OpenStudy (anonymous):

at the risk of repeating myself, is it \[2e^{\ln(x)+3}\] or \[2e^{\ln(x)}+3\]?

OpenStudy (anonymous):

I replied to you, but it's the second one. @satellite73

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

then start with \[2x+3=6\] and solve for \(x\) because \(e^{\ln(x)}=x\)

OpenStudy (anonymous):

so would I just have 2x = 3 and then divide the 3 by 2?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes (to your last post)

OpenStudy (raden):

agree ^^

OpenStudy (anonymous):

Thanks everyone! :) @byteme @RadEn @satellite73